12. A new temperature scale in o E is proposed wherein the boiling and freezing temperatures of water are

1250 o E and 85 o E respectively. If a fluid system has a temperature of 525 o E, determine its temperature

in a) o C and b) o R.


1
Expert's answer
2021-09-12T00:30:37-0400

Boiling temperature=1250°E

Freezing temperature=85°E

Boiling point-freezing point for this scale :

=(1250-85)°E=1165°E

Boiling point- freezing point for °C scale:

(100-0)°C=100°C

"\therefore" 100°C=1165°E

Hence, 1°C=11.65°E

For 525°E;

="525°E*\frac{1°C}{11.65°E}"

=45.06°C

°R=°C+273.15

=45.06+273.15

=318.21°R